# Models That Prove Their Own Correctness

### v1

@@ 4.3 Learning from annotated transcripts
However, when training Self-Proving model, it may be useful for it to first generate an “annotated” answer $\widetilde{a}$ which is then trimmed down to the actual answer $a$ to be sent to the verifier. We formally adapt the framework from Sections 3 and 4 to this setting in Appendix B, where we present Annotated Transcripts. This can be viewed as adding Chain-of-Thought (Wei et al., 2022) to the model. The Transcript Learning algorithm naturally extends to annotated transcripts as well.
@@ 5.1 Setup: Training transformers to predict the GCD of two integers
We choose the input distribution $\mu$ to be the log-uniform distribution on $[10^{4}]$, and train the transformer on sequences of the form $\mathbf{(x_{1}x_{2}y)}$, where $x_{1},x_{2}\sim\mu$ and $y=GCD(x_{1},x_{2})$. This is a scaling-down of Charton (2024), to allow single GPU training of Self-Proving transformers. In all of our experiments, we use a GPT model (Vaswani et al., 2017) with 6.3M parameters trained on a dataset of 1024K samples in batches of 1024. Full details are deferred to Appendix D.
@@
Following Charton (2024) as a baseline, we find that transformers can correctly compute the GCD with over $99\%$ probability over $(x_{1},x_{2})\sim\mu$. To what extent can they prove their answer? To answer this question, we first devise a natural proof system based on Bézout’s theorem. Its specification and formal guarantees are deferred to Appendix C. We denote its verification algorithm by $V$, and highlight some important features of the experimental setup:
@@
  • Completeness: For any $x_{1},x_{2},y\in[10^{4}]$ such that $y=GCD(x_{1},x_{2})$, there exists a proof $\mathbf{\pi}$ such that $V(\mathbf{x_{1}x_{2}y\pi})$ accepts. As detailed in Appendix C, the proof $\pi$ consists of a pair of integers who are Bézout coefficients for $x_{1},x_{2}$.
@@ 5.2 Models generalize beyond annotations
Figure 2: Verifiability with increasing amounts of annotation. $T$ is the number of steps added in Annotated Transcript Learning. Dashed lines bound the Verifiability of models that can only prove for integers up to a certain number of steps. Off chart to the left are bounds for depths 3 (47%), 4 (63%), and 5 (75%). Each $T$ was run with three seeds, with mean $\pm$ standard error depicted. See Appendix E for additional figures.
The proof $\mathbf{\pi}$ is annotated by including intermediate steps in its computation. Details are deferred to Appendix C; roughly speaking, we observe that the proof $\mathbf{\pi}$ for input $\mathbf{(a,b)}$ is obtained as the last element in a sequence $\mathbf{a,b,\pi_{1},\pi_{2},\dots}$ computed by the Euclidean algorithm. We annotate the proof $\mathbf{\pi}$ by prepending to it the sequence of Euclidean steps $\mathbf{(\pi_{1},\dots,\pi_{T})}$ up to some fixed cutoff $T$.
@@ 6 Conclusions
Acknowledgments and Disclosure of Funding
We are grateful to Micah Carroll for helpful comments. This research was supported by DARPA-TA1 under grant no. HR001119S0076, and by the Simons Collaboration on the Theory of Algorithmic Fairness.
@@ Appendix A Theoretical analyses for Section 4
In this section we provide a formal description and analysis of Transcript Learning (TL, Section 4.1) and Reinforcement Learning from Verifier Feedback (RLVF, Section 4.2). In Section A.1 we prove a convergence theorem for TL under convexity and Lipschitzness assumptions. Obtaining an analogous result for RLVF is more challenging; in lieu of a full analysis, we provide a lemma showing that the gradients estimated in the algorithm approximate the Verifiability of the model.
@@
We must first fully specify the theoretical framework in which our results reside. Continuing from Section 3, we define a learner as an algorithm $\Lambda$ with access to a family of autoregressive models $\{P_{\theta}\}_{\theta}$ and samples from the input distribution $x\sim\mu$. In our setting of Self-Proving models (and in consistence with the Interactive Proof literature), we give the learner the full specification of the verifier $V$. More formally,
@@
Throughout this section, we will refer to the transcript of an interaction between a verifier and a prover (see Figure 1). We will denote $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$, and for any index $s\in|\pi|$ we will write $\pi_{<s}\in\Sigma^{s-1}$ to denote the $s$-long prefix of $\pi$. Throughout this section, we will use $\pi\in\Sigma^{\ast}$ to denote the transcript of an interaction between a verifier and a prover.
@@ A.1 Transcript Learning
Fix a verifier $V$ in a proof system with $R\in\mathbb{N}$ rounds, where the verifier issues queries of length $L_{q}=|q_{i}|$ and the prover (model) responses with answers of length $L_{a}=|a_{i}|$. [^6] We can assume that queries (resp. answers) all have the same length by padding shorter ones. The query generator $V_{q}$ corresponding to $V$ takes as input a partial interaction and samples from the distribution over next queries by $V$. Formally, for any $r\leq R$, given input $x$, output $y$, and partial interaction $(q_{i},a_{i})_{i=1}^{r}$, $V_{q}(x,y,q_{1},a_{1},\dots,q_{r},a_{r})$ is a random variable over $\Sigma^{L_{q}}$. [^7] For completeness’ sake, we can say that when prompted with any sequence $z$ that does not encode an interaction, $V_{q}(z)$ is fully supported on a dummy token $\bot\in\Sigma$.
We assume that access to the verifier specification (Definition A.1) includes access to the query generator. After all, the verifier—who is assumed to be efficient—sampled from $V_{q}$ during the interaction. Moreover, we will assume that for any partial interaction and any sequence $q^{\prime}$, the learner is able to compute the probability that $q^{\prime}$ was the next query. In other words, we assume the learner can compute the probability density function of $V_{q}$.
@@
Fix a verifier $V$ in a proof system of $R\in\mathbb{N}$ rounds. A transcript generator $\mathcal{T}_{V}$ for $V$ is a randomized mapping from inputs $x\in\Sigma^{\ast}$ to transcripts $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})\in\Sigma^{\ast}$. For any input $x$, $\mathcal{T}_{V}(x)$ satisfies that for each $r\leq R$, the marginal of $\mathcal{T}_{V}(x)$ on the $r^{\text{th}}$ query ($q_{r}$) agrees with the corresponding marginal of the query generator $(V_{q})_{r}$.
A transcript generator $\mathcal{T}^{\ast}_{V}\coloneqq\mathcal{T}_{V}$ is honest if it is fully supported on transcripts $\pi^{\ast}$ for which the verifier accepts. [^8] WLOG we can assume that the prover sends her final answer $a_{R}$, the verifier’s decision is deterministic.
Notice that for any verifier $V$, there is a one-to-one correspondence between transcript generators and (possibly randomized) provers. We intentionally chose not to specify a prover in Definition A.3 to emphasize that transcripts can be “collected” independently of the honest prover (see completeness in Definition 3.2). As long as the generator is fully supported on honest transcripts, it can be used for Transcript Learning, described next (TL, Algorithm 1).
@@
$\theta_{i+1}\coloneqq\theta_{i}+\lambda\cdot\alpha_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}b_{r}(\theta_{i})a_{s}(\theta_{i})\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})$
@@
Convergence of TL is proven by a reduction to Stochastic Gradient Descent (SGD). Essentially, we are tasked with proving that TL estimates the gradient of the Verifiability of its model $P_{\theta}$. More precisely, TL estimates the gradient of a function that bounds the Verifiability from below. Maximizing this function therefore maximizes the Verifiability.
The lower-bounding function is the agreement of the transcript generator induced by $P_{\theta}$ with the provided honest transcript generator $\mathcal{T}^{\ast}_{V}$. More formally, we let $\mathcal{T}^{\theta}_{V}$ denote the transcript generator induced by the model $P_{\theta}$: for each $x$, $\mathcal{T}^{\theta}_{V}(x)$ is simply the distribution over transcripts of interactions between $V$ and $P_{\theta}$ on input $x$. We first prove that this function is differentiable.
Lemma A.4.
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. Fix an honest transcript generator $\mathcal{T}_{V}^{\ast}$. For any model $P_{\theta}$, it holds that
$\nabla_{\theta}\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right]=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\alpha_{0}(\theta)\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta)\cdot\alpha_{s}(\theta)\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)\right]$
@@
Proof of Lemma A.4.
Throughout this proof, expectations and probabilities will be over the same distributions as in the lemma statement. First,
$\Pr_{x,\pi^{*},\pi}\left[\pi=\pi^{*}\right]=\operatorname*{\mathbb{E}}_{x,\pi^{*}}\Pr_{\pi}[\pi=\pi^{*}],$
and so, by the linearity of the gradient,
$\nabla_{\theta}\Pr_{x,\pi^{*},\pi}\left[\pi=\pi^{*}\right]=\operatorname*{\mathbb{E}}_{x,\pi^{*}}\nabla_{\theta}\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right).$ (3)
The probability that the output of $V$ and $P_{\theta}$ on input $x$ is equal to a given transcript is (by the law of total probability) the product of probabilities that each of the tokens of the transcript is equal to the corresponding token of the given transcript, both tokens generated by $V$’s queries and by $P_{\theta}$’s answers, when conditioning on the prefix of the transcript.
Formally, consider any fixed $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$ and denote the random variable $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$. For any $r\in[R]$ denote the random variables $V_{q}^{<r}\coloneqq V_{q}(y,q_{1},a_{1},\dots,q_{r-1},a_{r-1})$ and $\mathcal{T}^{\theta,<r}_{V}\coloneqq\mathcal{T}^{\theta}_{V}(yq_{1}a_{1}\cdots a_{r-1}q_{r})$. Then,
@@
where Equation 4 uses independence of the verifier and models’ randomness, Equation 5 uses the autoregressive property of $P_{\theta}$ (Definition A.1), and Equation 6 is by definition of $\alpha_{s}$ and $\beta_{r}$.
@@
Let us focus on the last term. By Eq. 6,
@@
where Equation 8 is because $\log\beta_{r}(\theta)\coloneqq\log\Pr_{q^{\prime}\sim V_{q}(x\pi^{\ast}_{<t})}[q^{\prime}=q]$ is a constant and therefore has a gradient of zeros, and Equation 9 is by definition of $\vec{d}_{s}(\theta)$.
Combining Equations 6, 7 and 9 concludes the proof. ∎
@@
Theorem A.5.
[Theorem 4.1, formal] Fix a verifier $V$, an input distribution $\mu$, and an autoregressive model family $\{P_{\theta}\}_{\{\theta\in\mathbb{R}^{d}\}}$, and a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$. Fix an honest transcript generator $\mathcal{T}^{\ast}_{V}$ such that the expected agreement
$\mathrm{agree}_{\mathcal{T}^{\ast}_{V}}(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right]$
is convex in $\theta$. For any $\varepsilon>0$, we define
$B_{\varepsilon}$ $\coloneqq\min\left\{||\theta^{\ast}||:\mathrm{agree}_{\mathcal{T}^{\ast}_{V}}(\theta^{\ast})\geq 1-\varepsilon/2\right\}$
$\rho_{\varepsilon}$ $\coloneqq\max\left\{||\nabla_{\theta}\log p_{\theta}(z)||:z\in\Sigma^{\ast},||\theta||<B_{\varepsilon}\right\}$
For any $\varepsilon>0$ such that $B_{\varepsilon}$ and $\rho_{\varepsilon}$ are both finite, we denote by $\bar{\theta}$ the output of TL (Algorithm 1) running for number of iterations
$N=\frac{4R^{2}\cdot(L_{a}+1)^{2}\cdot\rho^{2}_{\varepsilon}\cdot B^{2}_{\varepsilon}}{\epsilon^{2}}$ (10)
and learning rate $\lambda=\varepsilon/2R^{2}L_{a}^{2}\rho$. Then the expected Verifiability (over the randomness of the samples collected by TL) is at least $1-\varepsilon$. That is,
@@
Our strategy is to cast TL as Stochastic Gradient Ascent (SGD). We follow Shalev-Shwartz and Ben-David [2014, Section 14.3], which is presented for Descent (SGD) but is equivalent up to sign change.
Let $\varepsilon$ such that $B_{\varepsilon}$ and $\rho_{\varepsilon}$ are finite be given. Since $B_{\varepsilon}<\infty$, let $\theta^{\ast}$ be such that $\mathrm{ver}_{V,\mu}(\theta^{\ast})\geq 1-\varepsilon/2$ and $||\theta^{\ast}||\leq B_{\varepsilon}$. To prove the theorem, it suffices to prove that
$\operatorname*{\mathbb{E}}[\mathrm{ver}_{V,\mu}(\bar{\theta})]\geq\mathrm{ver}_{V,\mu}(\theta^{\ast})-\varepsilon/2.$
Following the notation in Algorithm 1, for every iteration $i\in[N]$, $r\in[R]\cup\{0\}$ and $s\in S(r)$, the definition of $\rho_{\varepsilon}$ gives $||\vec{d}_{s}(\theta_{i})||\leq\rho_{\varepsilon}$. Thus, for each $i\in[N]$, we can bound the norm of the update step by
$\left\lVert a_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}b_{r}(\theta_{i})a_{s}(\theta_{i})\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert\\ =\left|a_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}b_{r}(\theta_{i})a_{s}(\theta_{i})\right)\right|\cdot\left\lVert\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert\\ \leq(R+1)\cdot L_{a}+\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\left\lVert\vec{d}_{s}(\theta_{i})\right\rVert\leq(R+1)\cdot(L_{a}+\rho_{\varepsilon}).$
For the above, we used the fact that $\alpha_{s}(\theta_{i}),\beta_{r}(\theta_{i})\leq 1$, the definition of the answer length $L_{a}$ ($|S(r)|=L_{a}$), and the triangle inequality on |||||$\cdot$||| | ⋅ | |. Therefore, by [Shalev-Shwartz and Ben-David, 2014, Theorem 14.8] and Lemma A.4, TL (Algorithm 1) satisfies
$\operatorname*{\mathbb{E}}_{\bar{\theta}}\left[\mathrm{agree}_{\mathcal{T}^{\ast}_{V}}\left(\bar{\theta}\right)\right]\geq\mathrm{agree}(\theta^{\ast})-\varepsilon/2\geq 1-\varepsilon,$
where the right inequality is by the choice of $\theta^{\ast}$. The proof follows by observing that, for any $\bar{\theta}$ (and in particular an expected one), it holds that $\mathrm{agree}_{\mathcal{T}^{\ast}_{V}}(\bar{\theta})\leq\mathrm{ver}_{V,\mu}(\bar{\theta})$; this is because, for any $x$, whenever the transcript generated by $\mathcal{T}^{\theta}(x)$ agrees with $\pi^{\ast}$, then the verifier accepts (by definition of $\pi^{\ast}$).

### v2

@@ 4.3 Learning from annotated transcripts
However, when training Self-Proving model, it may be useful for it to first generate an “annotated” answer $\widetilde{a}$ which is then trimmed down to the actual answer $a$ to be sent to the verifier. We adapt Sections 3 and 4 to this setting in Appendix C, where we present Annotated Transcripts. This can be viewed as adding Chain-of-Thought (Wei et al., 2022) to the model. The Transcript Learning algorithm naturally extends to annotated transcripts as well.
@@ 5.1 Setup: Training transformers to predict the GCD of two integers
We choose the input distribution $\mu$ to be the log-uniform distribution on $[10^{4}]$, and train the transformer on sequences of the form $\mathbf{(x_{1}x_{2}y)}$, where $x_{1},x_{2}\sim\mu$ and $y=GCD(x_{1},x_{2})$. This is a scaling-down of Charton (2024), to allow single GPU training of Self-Proving transformers. In all of our experiments, we use a GPT model (Vaswani et al., 2017) with 6.3M parameters trained on a dataset of 1024K samples in batches of 1024. Full details are deferred to Appendix E.
@@
Following Charton (2024) as a baseline, we find that transformers can correctly compute the GCD with over $99\%$ probability over $(x_{1},x_{2})\sim\mu$. To what extent can they prove their answer? To answer this question, we first devise a natural proof system based on Bézout’s theorem. Its specification and formal guarantees are deferred to Appendix D. We denote its verification algorithm by $V$, and highlight some important features of the experimental setup:
@@
  • Completeness: For any $x_{1},x_{2},y\in[10^{4}]$ such that $y=GCD(x_{1},x_{2})$, there exists a proof $\mathbf{\pi}$ such that $V(\mathbf{x_{1}x_{2}y\pi})$ accepts. As detailed in Appendix D, the proof $\pi$ consists of a pair of integers who are Bézout coefficients for $x_{1},x_{2}$.
@@ 5.2 Models generalize beyond annotations
Figure 2: Verifiability with increasing amounts of annotation. $T$ is the number of steps added in Annotated Transcript Learning. Dashed lines indicate Euclidean depth, that bound the Verifiability of models that prove only for integers up to a certain number of steps. Each $T$ was run with three seeds, with mean $\pm$ standard error depicted.
The proof $\mathbf{\pi}$ is annotated by including intermediate steps in its computation. Details are deferred to Appendix D; roughly speaking, we observe that the proof $\mathbf{\pi}$ for input $\mathbf{(a,b)}$ is obtained as the last element in a sequence $\mathbf{a,b,\pi_{1},\pi_{2},\dots}$ computed by the Euclidean algorithm. We annotate the proof $\mathbf{\pi}$ by prepending to it the sequence of Euclidean steps $\mathbf{(\pi_{1},\dots,\pi_{T})}$ up to some fixed cutoff $T$.
@@ 6 Conclusions
Acknowledgments
We are grateful to Micah Carroll and Avishay Tal for their helpful comments. This research was supported by DARPA-TA1 under grant no. HR001119S0076, and by the Simons Collaboration on the Theory of Algorithmic Fairness.
@@ Appendix A Theoretical analyses for Section 4
In this section we provide a formal description and analysis of Transcript Learning (TL, Section 4.1) and Reinforcement Learning from Verifier Feedback (RLVF, Section 4.2). In Section A.1 we prove a convergence theorem for TL under convexity and Lipschitzness assumptions. Obtaining an analogous result for RLVF is more challenging; in lieu of a full analysis, we provide a lemma showing that the gradients estimated in the algorithm approximate the Verifiability of the model in Section A.2.
@@
We must first fully specify the theoretical framework in which our results reside. Continuing from Section 3, we define a learner as an algorithm $\Lambda$ with access to a family of autoregressive models $\{P_{\theta}\}_{\theta}$ and samples from the input distribution $x\sim\mu$. In our setting of Self-Proving models (and in consistence with the Interactive Proofs literature), we give the learner the full specification of the verifier $V$. More formally,
@@
Throughout this section, we will refer to the transcript of an interaction between a verifier and a prover (see Figure 1). We will this transcript by $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$, and for any index $s\in|\pi|$ we will write $\pi_{<s}\in\Sigma^{s-1}$ to denote the $s$-long prefix of $\pi$.
@@ A.1 Transcript Learning
Fix a verifier $V$ in a proof system with $R\in\mathbb{N}$ rounds, where the verifier issues queries of length $L_{q}=|q_{i}|$ and the prover (model) responses with answers of length $L_{a}=|a_{i}|$. [^7] We can assume that queries (resp. answers) all have the same length by padding shorter ones. The query generator $V_{q}$ corresponding to $V$ takes as input a partial interaction and samples from the distribution over next queries by $V$. Formally, for any $r\leq R$, given input $x$, output $y$, and partial interaction $(q_{i},a_{i})_{i=1}^{r}$, $V_{q}(x,y,q_{1},a_{1},\dots,q_{r},a_{r})$ is a random variable over $\Sigma^{L_{q}}$. [^8] For completeness’ sake, we can say that when prompted with any sequence $z$ that does not encode an interaction, $V_{q}(z)$ is fully supported on a dummy sequence $\bot\cdots\bot\in\Sigma^{L_{q}}$.
We assume that access to the verifier specification (Definition A.1) includes access to the query generator. After all, the verifier—who is assumed to be efficient—samples from $V_{q}$ during the interaction. Moreover, we will assume that for any partial interaction and any sequence $q^{\prime}$, the learner is able to compute the probability that $q^{\prime}$ was the next query. In other words, we assume the learner can compute the probability density function of $V_{q}$.
@@
Fix a verifier $V$ in a proof system of $R\in\mathbb{N}$ rounds. A transcript generator $\mathcal{T}_{V}$ for $V$ is a randomized mapping from inputs $x\in\Sigma^{\ast}$ to transcripts $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})\in\Sigma^{\ast}$. For any input $x$, $\mathcal{T}_{V}(x)$ satisfies that for each $r\leq R$, the marginal of $\mathcal{T}_{V}(x)$ on the $r^{\text{th}}$ query $q_{r}$ agrees with the corresponding marginal of the query generator $(V_{q})_{r}$.
A transcript generator $\mathcal{T}^{\ast}_{V}\coloneqq\mathcal{T}_{V}$ is honest if it is fully supported on transcripts $\pi^{\ast}$ for which the verifier accepts.
Notice that for any verifier $V$, there is a one-to-one correspondence between transcript generators and (possibly randomized) provers. We intentionally chose not to specify a prover in Definition A.3 to emphasize that transcripts can be “collected” independently of the honest prover (see completeness in Definition 3.2). As long as the generator is fully supported on honest transcripts, it can be used for Transcript Learning (Algorithm 1 described below).
@@
$\theta_{i+1}\coloneqq\theta_{i}+\lambda\cdot\alpha_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta_{i})\alpha_{s}(\theta_{i})\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i}).$
@@
Convergence of TL is proven by a reduction to Stochastic Gradient Descent (SGD). Essentially, we are tasked with proving that TL estimates a surrogate of the Verifiability-gradient of its model $P_{\theta}$. More precisely, TL estimates the gradient of a function that bounds the Verifiability from below. Maximizing this function therefore maximizes the Verifiability.
The lower-bounding function is the agreement of the transcript generator induced by $P_{\theta}$ with the provided honest transcript generator $\mathcal{T}^{\ast}_{V}$. More formally, we let $\mathcal{T}^{\theta}_{V}$ denote the transcript generator induced by the model $P_{\theta}$: for each $x$, $\mathcal{T}^{\theta}_{V}(x)$ is simply the distribution over transcripts of interactions between $V$ and $P_{\theta}$ on input $x$. We first prove TL correctly estimates the gradient of $A(\theta)$ in its update step.
Lemma A.4 (TL gradient estimation).
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. Fix an honest transcript generator $\mathcal{T}_{V}^{\ast}$. Let $\theta$ be the parameters of a model $P_{\theta}$ such that
$A(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right]$
is differentiable in $\theta$. Then
$\nabla A(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ \pi^{\ast}\sim\mathcal{T}^{\ast}_{V}\end{subarray}}\left[\alpha_{0}(\theta)\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta)\cdot\alpha_{s}(\theta)\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta)\right]$
@@
Proof.
Throughout this proof, expectations and probabilities will be over the same distributions as in the lemma statement. First, by the law of total probability, and linearity of the gradient,
$\nabla A(\theta)\coloneqq\nabla_{\theta}\left(\Pr_{x,\pi^{*},\pi}\left[\pi=\pi^{*}\right]\right)=\nabla_{\theta}\left(\operatorname*{\mathbb{E}}_{x,\pi^{*}}\left[\Pr_{\pi}\left[\pi=\pi^{*}\right]\right]\right)=\operatorname*{\mathbb{E}}_{x,\pi^{*}}\left[\nabla_{\theta}\left(\Pr_{\pi}\left[\pi=\pi^{*}\right]\right)\right].$
Next, we use the law of total probability together with the autoregressive property of $P_{\theta}$ (Section 4) to switch from probabilities on transcripts, to products of next-token probabilities. Formally, consider any fixed input $x$, honest transcript $\pi^{\ast}=(y^{\ast},q_{1}^{\ast},a_{1}^{\ast},\dots,q_{R}^{\ast},a_{R}^{\ast})$, and denote a random transcript sampled from $\mathcal{T}_{V}^{\theta}(x)$ by $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$. For any $r\in[R]$ denote the random variables $V_{q}^{<r}\coloneqq V_{q}(y,q_{1},a_{1},\dots,q_{r-1},a_{r-1})$ and $\mathcal{T}^{\theta,<r}_{V}\coloneqq\mathcal{T}^{\theta}_{V}(yq_{1}a_{1}\cdots a_{r-1}q_{r})$. Then,
@@
where Equation 3 uses independence of the verifier and model’s randomness, Equation 4 uses the autoregressive property of $P_{\theta}$ (Definition A.1), and Equation 5 is by definition of $\alpha_{s}$ and $\beta_{r}$.
@@
Let us focus on the rightmost factor. By Equation 5,
@@
where Equation 7 is because $\log\beta_{r}(\theta)\coloneqq\log\Pr_{q^{\prime}\sim V_{q}(x\pi^{\ast}_{<t})}[q^{\prime}=q]$ is a constant and therefore has a gradient of zeros, and Equation 8 is by definition of $\vec{d}_{s}(\theta)$. Combining Equations 5, 6 and 8 concludes the proof. ∎
@@
Theorem A.5 (Theorem 4.1, formal).
Fix a verifier $V$, an input distribution $\mu$, and an autoregressive model family $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$, and a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$. Fix an honest transcript generator $\mathcal{T}^{\ast}_{V}$, and assume that the agreement function
$A(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ \pi^{*}\sim\mathcal{T}^{\ast}_{V}(x)\\ \pi\sim\mathcal{T}^{\theta}_{V}(x)\end{subarray}}\left[\pi=\pi^{*}\right]$
is concave and differentiable in $\theta$. For any $\varepsilon>0$, let $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ be upper-bounds such that the following conditions hold.
  • There exists $\theta^{\ast}\in\mathbb{R}^{d}$ with $||\theta^{\ast}||<B_{\mathrm{Norm}}$ such that $A(\theta^{\ast})\geq 1-\varepsilon/2$.
  • For all $\theta$, the logits of $P_{\theta}$ are $B_{\mathrm{Lip}}$-Lipschitz in $\theta$. That is,
    $\sup_{\begin{subarray}{c}\theta\in\mathbb{R}^{d}\\ z\in\Sigma^{\ast}\end{subarray}}||\nabla_{\theta}\log p_{\theta}(z)||\leq B_{\mathrm{Lip}}.$
  • In the proof system defined by $V$, the total number of tokens (over all rounds) is at most $C$.
Denote by $\bar{\theta}$ the output of TL running for number of iterations
$N\geq 4\cdot C^{2}\cdot\frac{B_{\mathrm{Norm}}^{2}\cdot B_{\mathrm{Lip}}^{2}}{\varepsilon^{2}}$
and learning rate $\lambda=B_{\mathrm{Norm}}/CB_{\mathrm{Lip}}\sqrt{N}$. Then the expected Verifiability (over the randomness of the samples collected by TL) of $\bar{\theta}$ is at least $1-\varepsilon$. That is,
@@
Our strategy is to cast TL as Stochastic Gradient Ascent and apply B.2. Let $\varepsilon$, $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ as in the theorem statement be given. Let $\theta^{\ast}$ be such that $A(\theta^{\ast})\geq 1-\varepsilon/2$ and $||\theta^{\ast}||\leq B_{\mathrm{Norm}}$.
First, notice that
$\operatorname*{\mathbb{E}}_{\bar{\theta}}\left[\mathrm{ver}_{V,\mu}(\bar{\theta})\right]\geq\operatorname*{\mathbb{E}}_{\bar{\theta}}[A(\bar{\theta})],$
This is because, for any $x$ and model $P_{\theta}$, whenever the transcript generated by $\mathcal{T}^{\theta}(x)$ agrees with $\pi^{\ast}$, then the verifier accepts (because $\pi^{\ast}$ is honest). Therefore, to prove the theorem it suffices to show that
$\operatorname*{\mathbb{E}}_{\bar{\theta}}[A(\bar{\theta})]\geq 1-\varepsilon.$
Following the notation in Algorithm 1, in every iteration $i\in[N]$ the norm of the update step is
$\left\lVert\alpha_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta_{i})\alpha_{s}(\theta_{i})\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert\\ =\left|\alpha_{0}(\theta_{i})\cdot\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta_{i})\alpha_{s}(\theta_{i})\right|\cdot\left\lVert\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i})\right\rVert\\ \leq 1\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\left\lVert\vec{d}_{s}(\theta_{i})\right\rVert,$
where the inequality is because $\alpha_{s}(\theta_{i})$ and $\beta_{r}(\theta_{i})$ are probabilities, so $\leq 1$. Continuing, we have
$\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\left\lVert\vec{d}_{s}(\theta_{i})\right\rVert\leq\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}B_{\mathrm{Lip}}\leq C\cdot B_{\mathrm{Lip}}.$
The first inequality is by definition of $B_{\mathrm{Lip}}$ as an upper-bound on the gradient of $P_{\theta}$’s logits. The second is because, by definition, $C$ is an upper bound on the number of tokens sent by the prover in the proof system, which is exactly the number of terms in the sum: $r$ indexes rounds, and $s$ indexes tokens sent in each round.
To conclude, Lemma A.4 shows that TL samples from a gradient estimator for $A(\theta)$, while the above equation shows that the gradient is upper-bounded by $C\cdot B_{\mathrm{Lip}}$. We can therefore apply B.2 to obtain
$\operatorname*{\mathbb{E}}_{\bar{\theta}}\left[A\left(\bar{\theta}\right)\right]\geq A(\theta^{\ast})-\varepsilon/2\geq(1-\varepsilon/2)-\varepsilon/2=1-\varepsilon,$
where the inequality is by definition of $\theta^{\ast}$.